OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)? 1 2 -2 -1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@UnkleRhaukus

jimthompson5910 (jim_thompson5910):

Subtract the x and y coordinates (in the same order) Then divide the y difference by the x difference to get the slope

OpenStudy (anonymous):

\[slope = \frac{ Y2- Y1 }{ X2 - X1 }\]

OpenStudy (anonymous):

the slope equation \[m=\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (anonymous):

i did it and didnt get an accurate answer

jimthompson5910 (jim_thompson5910):

What do you get when you subtract the y coordinates (y2 - y1) ?

OpenStudy (anonymous):

\[\frac{ -2 - 4 }{ -1 - (-4) }= \frac{ -6 }{ 3 }\]

OpenStudy (anonymous):

so its -2?

OpenStudy (anonymous):

sorry @jim_thompson5910

OpenStudy (anonymous):

ok thx @adel_demo and @jim_thompson5910 can you help me with one more question?

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

What type of line is the graph of 4x + y = -3? vertical horizontal rising falling

OpenStudy (anonymous):

this equation can be written as y=-4x-3 right @aleisha96

OpenStudy (anonymous):

@esamalaa yes

OpenStudy (anonymous):

this equation is similar to y=mx+c

OpenStudy (anonymous):

then the type of graph is ....?

OpenStudy (anonymous):

@esamalaa vertical?

OpenStudy (anonymous):

no :( can you draw it ?

OpenStudy (anonymous):

@esamalaa no it might be falling

OpenStudy (anonymous):

isnt this falling?

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

yes falling

OpenStudy (anonymous):

thanks @jim_thompson5910 @adel_demo @esamalaa

OpenStudy (anonymous):

@esamalaa @jim_thompson5910 another question What type of line is the graph of y = 7? vertical horizontal rising falling

OpenStudy (anonymous):

horizontal

OpenStudy (anonymous):



OpenStudy (anonymous):

Select the correct inequality for the graph below y < 3x - 1 y ≤ 3x - 1 y ≥ 3x - 1 y > 3x - 1

OpenStudy (anonymous):

@adel_demo @esamalaa @jim_thompson5910